### v1

# See What You Seek: Semantic Contextual Integration for Cloth-Changing Person Re-Identification

Abstract
Cloth-changing person re-identification (CC-ReID) aims to match individuals across multiple surveillance cameras despite variations in clothing. Existing methods typically focus on mitigating the effects of clothing changes or enhancing ID-relevant features but often struggle to capture complex semantic information. In this paper, we propose a novel prompt learning framework, Semantic Contextual Integration (SCI), for CC-ReID, which leverages the visual-text representation capabilities of CLIP to minimize the impact of clothing changes and enhance ID-relevant features. Specifically, we introduce Semantic Separation Enhancement (SSE) module, which uses dual learnable text tokens to separately capture confounding and clothing-related semantic information, effectively isolating ID-relevant features from distracting clothing semantics. Additionally, we develop a Semantic-Guided Interaction Module (SIM) that uses orthogonalized text features to guide visual representations, sharpening the model’s focus on distinctive ID characteristics. This integration enhances the model’s discriminative power and enriches the visual context with high-dimensional semantic insights. Extensive experiments on three CC-ReID datasets demonstrate that our method outperforms state-of-the-art techniques. The code will be released at github.
Index Terms:
Person Re-Identification, Clothing Changes, Vision-Language Models, Prompt Learning, Semantic Integration.I Introduction
Person re-identification (Re-ID) is a critical task that involves identifying individuals across different locations over time, with significant applications in video surveillance and smart city infrastructure [1, 2, 3]. Traditional Re-ID methods [4, 5, 6, 7] focus on clothing features, such as texture and color, to improve performance. However, these methods become ineffective when individuals change clothes, highlighting the need for robust approaches that can handle clothing variations in real-world settings.
To address the challenge of clothing changes in Re-ID (CC-ReID), extensive research has yielded promising results [8, 9]. Existing methods can be broadly categorized into two categories: those aiming to reduce the influence of clothing and those focusing on enhancing ID-relevant information. The first category seeks to minimize the impact of clothing features, as illustrated in Fig. 1(a). For instance, CAL [10] uses adversarial learning to reduce reliance on clothing by penalizing predictions based on the RGB modality, while AIM [11] employs causal intervention with a dual-branch model to mitigate clothing bias. The second category emphasizes ID-relevant features using auxiliary cues such as pose, gait, or human parsing networks. For example, GI-ReID [12] uses gait information to learn clothing-agnostic representations, and SCNet [13] leverages human body parsing network to learn identity-relevant features through semantic consistency constraints.
Although these methods address clothing changes, they typically focus either on minimizing clothing influence or enhancing ID cues through explicit features like body contours. We argue that implicit positive factors such as hair, glasses, or backpacks are embedded in the visual features but are challenging to define explicitly or separately. As shown in Fig. 1(b), our goal is to remove negative influences while making implicit positive factors more explicit.
Recent advancements in visual-language learning have shown great potential in enhancing visual recognition and understanding. Contrastive Language-Image Pre-training (CLIP) [14], a prominent cross-modal pre-training model, has effectively bridged the gap between visual data and natural language, enabling more context-aware and interpretatively rich analysis. CLIP has achieved remarkable success in various downstream visual tasks [15, 16]. In Re-ID, recent work [17, 18] demonstrates that linking visual content with corresponding language descriptions allows models to capture complex semantics associated with individuals. However, these studies have primarily focused on traditional Re-ID, applying CLIP to CC-ReID has not yet been explored. This motivates us to leverage prompt learning to reveal high-dimensional semantic information, thereby enhancing visual-textual connections and making implicit positive factors more explicit.
In this paper, we propose the Semantic Contextual Integration (SCI) framework for CC-ReID, which utilizes CLIP to capture high-dimensional semantic features. Since CLIP struggles to selectively ignore aspects like clothing through negative prompts, we introduce a Semantic Separation Enhancement (SSE) module to minimize the influence of clothing by isolating semantic-level negative factors and preserving key positive features for CC-ReID.
Specifically, we use dual learnable text tokens to separately capture confounding and clothing-related semantic information. We then apply an orthogonalization process on two distinct sets of text features to filter out clothing-related features from the confounding ones. SSE not only reduces the negative impact of clothing but also retains important positive semantic factors. These refined text features are used to guide the visual encoder.
To further emphasize implicit positive factors in visual images, we introduce a Semantic-Guided Interaction Module (SIM), which directs the visual encoder’s extraction process using the filtered text features. This allows the model to focus on ID-relevant elements beyond explicit features such as body contours. By leveraging CLIP’s text-visual integration capabilities, our method enhances the interaction between the two modalities, improving visual feature refinement and multi-modal alignment.
Our contributions are summarized fourfold:
We propose the Semantic Contextual Integration (SCI) framework to leverage semantic information in CC-ReID, removing negative factors, emphasizing implicit positive elements, and refining visual representations.
We introduce the Semantic Separation Enhancement (SSE) module to filter and refine text-level features, improving the model’s ability to isolate key semantic information.
We design the Semantic-Guided Interaction Module (SIM) to guide visual representations using refined text features, enhancing multi-modal integration and alignment.
Extensive experiments on three standard CC-ReID datasets demonstrate the effectiveness of our approach, providing a robust solution to real-world Re-ID challenges.
II Related Work
II-A Cloth-Consistent Person Re-ID
Person Re-ID under consistent clothing conditions has been extensively studied [19, 20, 21]. These methods primarily rely on clothing color and texture to extract discriminative features. Feature representation learning in this context can be categorized into three types: global features, local features, and auxiliary information-based features.
Global feature representation methods extract a single feature vector for each person image [22]. Local feature representation methods aggregate part-based features to address misalignments, using techniques like human parsing [23] for semantic part detection, horizontal division [24] for structured feature aggregation, hard region-level and soft pixel-level attention [25] for optimized features, and IANet [26] for enhanced small visual cues. Auxiliary information-based methods incorporate additional data, such as semantic attributes [19] or generated/augmented samples [20], to enhance generalization by embedding richer contextual information. And ISP [27] locates human body parts and personal belongings at pixel-level for aligned with labels.
Although effective in handling variations in appearance, background, and camera perspectives, these methods assume consistent clothing, leading to performance declines in scenarios with clothing changes, which are common in real-world applications. This limitation has spurred increased focus on addressing cloth-changing person Re-ID challenges.
II-B Cloth-Changing Person Re-ID
To advance CC-ReID, several datasets such as PRCC [28], LTCC [29], Celeb-reID [30], and VC-Clothes [31] have been introduced, providing diverse data for effective model evaluation. The core challenge in CC-ReID is to learn features that reliably identify individuals despite clothing variations.
There are three primary approaches to this challenge. The first leverages auxiliary soft-biometric information. For example, CAMC [32] incorporates body shape semantics into ID-related features, enhancing robustness to clothing changes, while FSAM [33] learns fine-grained body shape features to complement clothing-independent features. M2Net [34] utilizes contour and human parsing information to learn features that are robust to appearance changes.
The second approach involves disentangling features to isolate clothing-irrelevant information and reduce clothing bias. For instance, CAL [10] uses clothes-based adversarial loss to emphasize clothing-irrelevant features, and AIM [11] applies causal intervention to automatically eliminate clothing effects.
The third approach employs data augmentation or adaptation techniques. Pos-Neg [8] and CCFA [9] enhance clothing color and texture diversity without additional data collection, improving robustness to clothing variations. RCSANet [35] explicitly constructs a clothing status, which is used to enhance the robustness of ID features for handling both cloth-changing and no-cloth-changing cases.
Despite these advancements, existing methods primarily rely on visual inputs for model training, limiting their capacity for deep semantic understanding. In this paper, we propose a novel method that integrates visual and natural language information, utilizing CLIP to enhance the interaction of the visual content with corresponding textual descriptions.
II-C Vision-Language Learning
Vision-language pre-training (VLP) has recently shown significant promise in enhancing the performance of various downstream vision tasks by training models to associate images with textual information. CLIP [14], a leading VLP model, employs paired image and text encoders to bridge the gap between natural language and visual content. By using similarity learning, CLIP aligns textual and visual representations in a shared space, enhancing tasks such as image captioning and classification.
Building on CLIP, recent advances in prompt learning, such as CoOp [36], transform prompt context words into learnable vectors. CoCoOp [37] extends this by generating input-conditional tokens for each image through a lightweight network. DenseCLIP [38] adapts this approach for dense prediction models, introducing pixel-text matching to utilize contextual information from a pre-trained language model.
In person Re-ID, CLIP-ReID [17] is the first to use ID-specific learnable tokens and a two-stage training strategy to fully utilize text for specifying visual concepts. CCLNet [18] advances this approach with a novel prompt learning paradigm for unsupervised visible-infrared Re-ID. However, these studies primarily focus on scenarios with consistent clothing, leaving CLIP’s potential in CC-ReID largely unexplored. In this paper, we propose the CLIP-CCReID framework to mitigate the impact of clothing changes in CC-ReID.
III Proposed Method
III-A Preliminaries and Overview
The large-scale pre-trained vision-language model CLIP [14] demonstrates significant advantages in learning joint representations of images and text. CLIP consists of two encoders: a visual encoder and a text encoder . CLIP aligns the embedding spaces of visual and textual data during pre-training using a contrastive objective.
For downstream classification tasks, a simple and effective method [36] constructs text prompts using a template like “a photo of a [CLS].”, where [CLS] is replaced with specific class names. Given an image, CLIP computes similarities between the image embedding and the text prompt embeddings, the process involves transforming the image and text into a shared high-dimensional space where their relationships can be quantitatively assessed. CLIP then selects the class corresponding to the highest similarity as the prediction.
Building upon CLIP, CoOp [36] introduces learnable textual contexts, represented as “ [CLS].”, where () are learnable tokens. This approach enhances the model’s transferability and performance. Leveraging the capabilities of CLIP and CoOp, we employ these textual descriptions to achieve deeper semantic understanding for CC-ReID, as illustrated in Fig. 2, enabling more robust and accurate modeling despite clothing changes.
III-B Semantic Separation Enhancement
Since CLIP cannot learn negative semantics or ignore specific elements like clothing through negative prompts, we propose the Semantic Separation Enhancement (SSE) module to address this limitation. This module mitigates the impact of clothing variations by subtracting negative semantic elements, preserving only the high-dimensional positive features crucial for CC-ReID. The complete algorithm is summarized in Algorithm 1.
We define dual textual prompts as follows:
where () are learnable tokens initialized randomly, and denotes the number of tokens. In the first training stage, we use the pre-trained visual encoder and text encoder to extract image and dual text features, freezing the encoders’ parameters to focus on optimizing the text tokens . This allows us to learn contextual representations by updating the text tokens, thereby acquiring distinct textual representations for each ID and their clothing:
where and represent the ID-specific text features and clothing text features, respectively.
After extracting the dual text representations, we project onto to reduce the influence of clothing and emphasize the implicit positive factors. This step clarifies the relation between clothing representations and ID representations, thereby minimizing the impact of clothing variations. The computation process is as follows:
We then perform orthogonalization by subtracting the projection from , removing the component in the direction of at the textual level. This process ensures that the features are aligned with the positive features while removing the influence of negative factors. The orthogonalization process is:
where is the feature orthogonalized to align with and orthogonal to .
To optimize the text features, we design a loss function that maximizes the similarity between and , while minimizing the similarity with . This enables the model to better capture positive factors without being affected by negative aspects, such as clothing changes. The similarity calculations are:
where denotes the average value, and represents the cosine similarity function. The loss function is as follows:
where and are weighting factors controlling the contributions of ID and clothing similarities.
Finally, we calculate the image-to-text contrastive loss :
where and are paired visual and text embeddings, denotes the similarity function, and is the batch size. Since multiple images in a batch may belong to the same ID, meaning there may be multiple positive samples, the text-to-image contrastive loss is calculated as:
where is the set of all positive indices for in the batch, and is the text embedding corresponding to label . Therefore, the overall loss function is:
III-C Semantic-Guide Interaction Module
Previous work [10, 17, 39] shows that reducing clothing bias in visual or textual features improves performance in cloth-changing scenarios. However, these methods typically treat visual and textual branches independently, lacking interaction. Our method refines visual representations by leveraging clothing-irrelevant textual features, enhancing the interaction between visual and textual branches. This integration introduces more robust and invariant descriptors into the visual processing, improving the model’s ability to identify individuals across clothing changes. The algorithm is summarized in Algorithm 2.
As illustrated in Fig. 3, we employ a cross-attention mechanism in a Transformer decoder [40] to model interactions between visual and textual data. Instead of traditional self-attention, we use a non-local operation [41] to capture global context, ensuring the model can integrate information from distant but relevant positive parts of the visual image, which is crucial in cloth-changing scenarios. During this stage, we update the visual encoder while keeping other components frozen. The computation is as follows:
where is the input feature map, , , and are linear transformations, and is the number of elements in the feature map. captures global context information.
The final visual feature map is:
where contains learnable parameters. is the final output feature map obtained by adding this context-aware feature to the original feature map.
We further enhance interaction by using text embeddings as queries and visual embeddings as keys and values. This cross-attention mechanism refines visual features using text information, ensuring alignment with relevant semantic context:
where is the dimensionality of the keys. This implementation strategically encourages the text features to find the most relevant visual clues.
Finally, we update the visual features. This process ensures that the visual features are enriched with relevant semantic information from the text, leading to more robust and contextually aware visual representations:
where is a learnable parameter controlling the degree of refinement applied to the visual features.
This module employs an interaction-to-alignment approach, ensuring that text embeddings guide the visual encoding process, transforming implicit information into explicit visual cues. By emphasizing positive factors and reducing negative ones, this method enhances the model’s ability to consistently identify individuals across various clothing scenarios.
For the objective function, we incorporate cross-entropy loss and clothes-based adversarial loss [10] to optimize the visual encoder:
where is the true label for the -th sample, and is the predicted probability of the true label . The clothes-based adversarial loss is expressed as:
where is the number of clothing categories, is the clothes classifier, is the weight for the -th class, is the feature of sample , and is a temperature parameter.
To fully leverage the capabilities of CLIP, we calculate the image-to-text cross-entropy loss as:
where label smoothing is applied to . The total loss for the SIM module is:
IV Experimental Results
IV-A Datasets and Evaluation Metrics
We evaluate our proposed method on three standard cloth-changing datasets: LTCC [29], PRCC [28], and VC-Clothes [31], following CAL [10] and AIM [11]. Additionally, we also evaluate our method on two traditional Re-ID datasets: Market-1501 [42], and MSMT17 [43], to validate its applicability.
LTCC dataset contains 17,138 images of 152 IDs captured by 12 cameras. The training set comprises 9,576 images of 77 IDs, and the testing set includes 7,543 images (493 queries and 7,050 gallery images) of 75 IDs. This long-term Re-ID dataset features frequent clothing changes and varied environmental conditions, with each ID having between 2 and 14 outfits.
PRCC dataset consists of 33,698 images of 221 IDs captured by three cameras. The training set contains 22,898 images of 150 IDs, while the testing set includes 10,800 images of 71 IDs. Each person has two outfits, with cameras A and B capturing the same clothes and camera C capturing a different outfit.
VC-Clothes dataset is a synthetic dataset generated using the GTA5 game engine. It comprises 9,449 images of 256 IDs across both training and testing sets, featuring 1,241 clothing items across four camera views.
Market-1501 dataset includes 1,501 individuals recorded by six different cameras. It consists of 12,936 images of 751 IDs for training and 19,732 images of 750 IDs for testing. The evaluation is conducted using the single query setting.
MSMT17 dataset includes 1,041 individuals captured by 15 different cameras. The training set contains 32,621 images of 1,041 IDs, while the testing set includes 51,027 images of 11,659 IDs. This dataset is notably large and complex.
We use the cumulative matching characteristic (CMC) curve and mean average precision (mAP) as evaluation metrics. Rank- in CMC measures the likelihood of retrieving the correct ID in the top- results, while mAP captures the average retrieval performance across all queries.
Evaluation is conducted under three settings: 1) General setting: Excludes same-ID and same-camera samples, using both cloth-changing and clothing-consistent samples. 2) Same-clothes setting: Excludes same-ID and same-camera samples, using only clothing-consistent samples. 3) Cloth-changing setting: Excludes same-ID, same-camera, and same-clothing samples, using only cloth-changing samples.
IV-B Implementation Details
We adopt a modified ResNet-50 pre-trained on CLIP as the backbone for feature extraction. Images are resized to pixels, with a batch size of 64. Training employs the Adam optimizer [44] with random horizontal flipping, cropping, and erasing [45] for data augmentation. A global attention pooling layer reduces the feature dimension from 2048 to 1024, matching the text feature dimension, which is scaled from 512 to 1024.
In the first stage, two prompt learners are trained for 60 epochs on LTCC, VC-Clothes, Market-1501, and MSMT17, and 20 epochs on PRCC, with an initial learning rate of adjusted using a cosine decay schedule. In the second stage, the visual encoder is trained for 120 epochs on LTCC, VC-Clothes, Market-1501, and MSMT17, and 20 epochs on PRCC, with learning rates initially set to and reduced by a factor of 10 at the 40th and 70th epochs. All experiments are conducted on the PyTorch platform using a single A100 GPU.
IV-C Comparison with State-of-the-Art Methods
We compare our method with state-of-the-art techniques on LTCC, PRCC, and VC-Clothes, including traditional Re-ID methods HACNN [25], PCB [24], IANet [26], and ISP [27], and specialized CC-ReID methods FSAM [33], RCSANet [35], CAL [10], GI-ReID [12], M2Net [34], and AIM [11].
| Methods | Venue | LTCC | PRCC | VC-Clothes | |||||||||
| General | Cloth-changing | Same-clothes | Cloth-changing | General | Cloth-changing | ||||||||
| R-1 | mAP | R-1 | mAP | R-1 | mAP | R-1 | mAP | R-1 | mAP | R-1 | mAP | ||
| HACNN [25] | CVPR ’18 | 60.2 | 26.7 | 21.6 | 9.3 | 82.5 | 84.8 | 21.8 | 23.2 | 68.6 | 69.7 | 49.6 | 50.1 |
| PCB [24] | ECCV ’18 | 65.1 | 30.6 | 23.5 | 10.0 | 99.8 | 97.0 | 41.8 | 38.7 | 87.7 | 74.6 | 62.0 | 62.2 |
| IANet [26] | CVPR ’19 | 63.7 | 31.0 | 25.0 | 12.6 | 99.4 | 98.3 | 46.3 | 45.9 | - | - | - | - |
| ISP [27] | ECCV’20 | 66.3 | 29.6 | 27.8 | 11.9 | 92.8 | - | 36.6 | - | 94.5 | 94.7 | 72.0 | 72.1 |
| FSAM [33] | CVPR’21 | 73.2 | 35.4 | 38.5 | 16.2 | - | - | - | - | 94.7 | 94.8 | 78.6 | 78.9 |
| RCSANet [35] | ICCV’21 | - | - | - | - | 100.0 | 97.2 | 50.2 | 48.6 | - | - | - | - |
| CAL [10] | CVPR’22 | 73.4 | 39.4 | 38.0 | 17.2 | 100.0 | 99.6 | 54.7 | 55.4 | 93.1 | 88.3 | 82.6 | 81.7 |
| GI-ReID [12] | CVPR’22 | 63.2 | 29.4 | 23.7 | 10.4 | - | - | - | - | - | - | 64.5 | 57.8 |
| M2Net [34] | ACM MM’22 | - | - | - | - | 99.5 | 99.1 | 59.3 | 57.7 | - | - | - | - |
| AIM [11] | CVPR’23 | 75.7 | 41.3 | 41.8 | 17.9 | 100.0 | 99.7 | 56.3 | 56.5 | 91.2 | 84.9 | 81.0 | 75.7 |
| CLIP-ReID [17] | AAAI’23 | 73.0 | 36.5 | 38.3 | 15.9 | 100.0 | 99.5 | 57.0 | 55.3 | 93.1 | 86.9 | 85.9 | 79.3 |
| SCI (Ours) | – | 75.7 | 40.6 | 42.1 | 18.6 | 99.6 | 97.7 | 59.8 | 56.2 | 94.9 | 89.2 | 89.2 | 83.1 |
As shown in Table I, CC-ReID methods significantly outperform general Re-ID methods, primarily because general methods rely heavily on clothing for identification, which is less effective in long-term surveillance with clothing changes. Our SCI method achieves competitive results by effectively eliminating ID-irrelevant features and enhancing positive ID elements in visual images.
On LTCC, SCI surpasses the baseline [17] in the cloth-changing setting, with improvements of 3.8% in Rank-1 accuracy and 2.7% in mAP. Although our mAP is slightly lower than AIM [11] in the general setting, this is likely due to LTCC’s inherent challenges, where models often rely on clothing information. By filtering out such information, SCI experiences minor declines in non-cloth-changing scenarios, but these are negligible given the overall benefits. Notably, SCI achieves these results without incorporating additional modalities or modules, unlike other methods. The primary advantage of our method lies in its ability to avoid carrying clothing-relevant information, which is often unavoidable in methods relying on auxiliary data such as poses and gaits. Compared with non-auxiliary-based methods, SCI outperforms CAL by 12.7% in Rank-1 and 6.3% in mAP in the cloth-changing setting on VC-Clothes. On PRCC, SCI also surpasses AIM by 3.5% in Rank-1 accuracy for clothing changes. Although the differences on other metrics are less pronounced due to PRCC’s simpler structure (only two outfits per ID), SCI still shows notable improvements. On VC-Clothes, SCI demonstrates significant gains in the cloth-changing setting, with a 7.4% higher mAP than AIM, underscoring the effectiveness of our approach in challenging scenarios.
IV-D Ablation Studies and Analysis
We conduct ablation studies to assess the effectiveness of the components in SCI and analyze parameter impacts.
IV-D1 Effectiveness of Semantic Separation Enhancement
We first evaluate the baseline without additional modules, where the model relies solely on a prompt “a photo of a person” for classification. This approach shows reduced performance. Introducing the SSE module, which uses two prompts (“a photo of a person” and “a photo of the clothes”) to learn features separately at the textual level, improves Rank-1 and mAP by around 2.0% across datasets, demonstrating its ability to filter clothing features.
These results indicate: 1) The original generic textual representations are confounding, containing both ID and clothing features, leading the model to focus excessively on cloth-relevant aspects. 2) Separating these features at the textual level is simple yet effective. These findings demonstrate the effectiveness of setting specific prompts for each factor and the robustness of our SSE module against both cloth-changing and general scenarios.
| Baseline | SSE | SIM | LTCC | PRCC | VC-Clothes | |||||||||
| General | Cloth-changing | Same-clothes | Cloth-changing | General | Cloth-changing | |||||||||
| R-1 | mAP | R-1 | mAP | R-1 | mAP | R-1 | mAP | R-1 | mAP | R-1 | mAP | |||
| ● | ○ | ○ | 73.0 | 36.5 | 38.3 | 15.9 | 100.0 | 99.5 | 57.0 | 55.3 | 93.1 | 86.9 | 85.9 | 79.3 |
| ● | ● | ○ | 76.5 | 39.7 | 39.8 | 18.3 | 100.0 | 98.4 | 59.0 | 54.7 | 94.1 | 88.0 | 87.6 | 81.1 |
| ● | ○ | ● | 72.8 | 36.3 | 37.8 | 16.1 | 99.7 | 98.0 | 52.4 | 49.9 | 94.0 | 88.9 | 88.4 | 82.4 |
| ● | ● | ● | 75.7 | 40.6 | 42.1 | 18.6 | 99.6 | 97.7 | 59.8 | 56.2 | 94.9 | 89.2 | 89.2 | 83.1 |
IV-D2 Effectiveness of Semantic-Guided Interaction Module
We further assess SIM by integrating it with the baseline model. As shown in Table II, using SIM alone often degrades performance, particularly in the cloth-changing scenario on PRCC, where Rank-1 drops by 4.6% and mAP decreases from 55.3% to 49.9%.
The decline in performance in this set of ablation experiments is primarily due to the absence of the SSE module. Typically, the SSE module generates final textual features that guide the visual representations during the updating process of the visual encoder in SIM. However, with only a single prompt “a photo of a person” used to extract textual representations, many ID-irrelevant negative factors, particularly clothing factors, are included. This results in an excessive emphasis on clothing-relevant representations during guidance in SIM, reducing overall performance.
| (a) Baseline (stage1) | (b) Baseline (stage2) | (c) Ours (stage1) | (d) Ours (stage2) |
IV-D3 Feature Distribution Analysis
To illustrate SCI’s effectiveness, we employ t-SNE visualization [46], as depicted in Fig. 4. This visualization plots the statistical distribution for 20 randomly selected categories from LTCC, comparing the latent space distributions at different stages of the baseline model with our approach.
From Figs. 4(a) and (c), it is observable that both the baseline and our method exhibit disorganized feature distributions with blurred class boundaries. During this stage, only the prompts are trained while the visual and text encoders remain frozen, limiting their capability to effectively extract diverse representations. In contrast, as shown in Figs. 4(b) and (d), the feature distributions for both methods become clearer, indicating the effectiveness of the representations derived from the text-image similarity computations.
Notably, the clusters in Fig. 4(d) are more compact and distinct compared to those in Fig. 4(b). The red and green dashed circles in Fig. 4(b) enclose samples with considerable scatter, which are better clustered together in Fig. 4(d). Therefore, this visualization not only serves as compelling validation of our capability to effectively extract features and discriminate IDs but also highlights our method’s potential to significantly advance the field in cloth-changing scenarios.
IV-D4 Visualization of Feature Maps
To illustrate the distinctions between the baseline [17] and SCI, we present heatmaps of regions of interest in Fig. 5, demonstrating performance across LTCC, VC-Clothes, and PRCC. Fig. 6 shows VC-Clothes for the same ID but with different outfits, illustrating the regions of interest that our SCI model focuses on under clothing variations. These visualizations provide insights into the unique differences in feature activation patterns.
As shown in Fig. 5, the baseline method tends to focus on more dispersed areas, which include negative factors, potentially detracting from its effectiveness. Conversely, SCI emphasizes highly discriminative ID features, such as the person’s head, shoulders, lower body, and shoes. This specific emphasis on footwear aligns with findings from previous studies [47], which suggest that footwear often remains consistent across different outfits and scenarios. This indicates our method’s enhanced ability to capture crucial positive factors when encountering challenging scenarios.
In Fig. 6, for each person, the corresponding heatmaps illustrate how the feature extraction model highlights various regions of interest across different outfits, reflecting the model’s focus on certain body parts regardless of clothing changes. However, there are some limitations. As indicated by the red circles, the model tends to focus excessively on areas like the knees. If people wear long pants, this focus could potentially affect the results. We are aware of this issue and will continue to optimize our model to address it.
| (a) R-1 Accuracy on cloth-changing | (b) mAP Accuracy on cloth-changing | (c) R-1 Accuracy on general | (d) mAP Accuracy on general |
IV-D5 Comparative Analysis of Mechanisms in SIM
As shown in Fig. 7, we evaluate the effectiveness of the non-local operation [41] compared to self-attention mechanisms [40] in leveraging textual information obtained from the SSE module to guide visual representations. In cloth-changing scenarios, the non-local mechanism yields superior performance, achieving 42.1% in Rank-1 accuracy and 18.6% in mAP, as depicted in Figs. 7(a) and (b), thereby outperforming the self-attention mechanism by margins of 4.6% and 2.4%, respectively.
This underscores the non-local mechanism’s enhanced capability to capture comprehensive global context, which is vital for integrating information from disparate yet significant segments of the visual field in cloth-changing scenarios. In general scenarios, as illustrated in Figs. 7(c) and (d), both mechanisms exhibit comparable efficacy, with the non-local mechanism slightly edging out self-attention by a 2.0% increase in mAP. This suggests the non-local mechanism is robust and can be effectively deployed for person Re-ID tasks.
IV-D6 Loss Functions Analysis
We conduct experimental evaluations to explore the effectiveness of (9), as demonstrated in Fig. 8(a). This loss function significantly improves performance by optimizing for higher ID similarity and reducing clothing similarity, making it particularly effective for applications that need to distinguish personal ID from external features like clothing, relevant in tasks such as cloth-changing scenarios. It helps the model focus on positive factors while ignoring negative variations such as clothing changes, thereby enhancing adaptability and accuracy in complex environments.
| Methods | Venue | Market1501 | MSMT17 | |||
| R-1 | mAP | R-1 | mAP | |||
| Cloth- consistent | AGW [48] | TPAMI’21 | 95.1 | 87.8 | 68.3 | 49.3 |
| reID-NAS [49] | TNNLS’22 | 95.1 | 85.7 | 79.5 | 53.3 | |
| CLIP-ReID [17] | AAAI’23 | 95.7 | 89.8 | 84.4 | 63.0 | |
| IRM [50] | CVPR’24 | 96.5 | 93.5 | 86.9 | 72.4 | |
| Cloth- changing | CAL [10] | CVPR’22 | 85.0 | 67.8 | 65.1 | 38.4 |
| AIM [11] | CVPR’23 | 70.3 | 48.0 | 26.6 | 12.0 | |
| SCI (Ours) | – | 85.7 | 68.6 | 73.5 | 47.4 | |
IV-D7 Cloth-consistent Re-ID Scenario
To validate the applicability of our method in cloth-consistent scenarios, we conducted experiments using our cloth-changing approach on two traditional ReID datasets: Market1501 and MSMT17. For ease of comparison, we also incorporated cloth-consistent methods: AGW [48], reID-NAS [49], CLIP-ReID [17], and IRM [50]. This enables a comprehensive evaluation of the performance in varying scenarios.
As shown in Table III (cloth-consistent), existing Re-ID methods designed for traditional datasets have achieved progressively better performance, even on larger-scale datasets like MSMT17. While our method is specifically designed to address the more realistic scenario of clothing changes, we are also interested in evaluating its performance on conventional datasets. From Table III (cloth-changing), it can be observed that our SCI method achieves the best performance on both Market1501 and MSMT17 among cloth-changing methods. Although there is still a gap compared to cloth-consistent methods, this highlights an important direction for future research and improvement.
IV-D8 Visualization of Retrieved Examples
We present the visualization of the top-10 retrieved results from randomly selected query examples on LTCC in Fig. 9. It observes that the retrieval results are predominantly optimistic, reflecting a high level of accuracy and relevance in most cases. However, some incorrect results still persist, which could be due to challenges such as varying viewpoints, low lighting conditions, or occlusions within the images. These issues underscore potential areas for future research. Despite these challenges, the overall result strongly supports the effectiveness of our SCI, demonstrating its robustness across a variety of scenarios.
IV-D9 Parameter Selection
We evaluate the parameters and specified in (9). The results for LTCC are displayed in Fig. 8(b). We determine that and provide the best performance in most scenarios, optimizing the balance between enhancing ID and reducing clothing similarity, while feature orthogonalization minimizes interference between ID and clothing features. This setting allows the model to effectively discern between the positive representations and their negative factors, such as clothing. Additionally, and obtain the second-best results. This analysis underscores the effectiveness and adaptability of our proposed method in real-world surveillance applications.
IV-D10 Efficiency Analysis
To verify the efficiency and performance of SCI, we analyze several computational and training metrics across different methods, detailed in Table IV. AIM [11] utilizes a dual-branch ResNet model, achieving an R-1 accuracy of 41.8%, which surpasses CAL [10], another ResNet-based model. Although AIM requires higher GFLOPs compared to CAL, it significantly outperforms CAL in terms of accuracy with the same parameter count.
In CLIP-ReID [17] and our method, we integrate textual modality information, transitioning from a single-modality to a multimodal approach. This justifies the observed increases in both GFLOPs and parameters, which remain within acceptable limits. Our SCI achieves an R-1 accuracy of 42.1%, demonstrating the value of incorporating textual information. This allows the model to better distinguish between different IDs, especially in challenging scenarios. Overall, our method provides a good balance between computational cost and performance enhancements.
V Conclusion and Limitation
In this study, we propose the Semantic Contextual Integration (SCI) network, a novel method incorporates prompt learning to address cloth-changing person re-identification (CC-ReID). The SCI network includes a Semantic Separation Enhancement (SSE) module that generates ID-specific and clothing prompts to effectively isolate clothing semantics (negative factors) from ID information, thereby preserving high-dimensional semantic details (positive factors) embedded in visual images. Additionally, the Semantic-Guided Interaction Module (SIM) leverages textual cues to guide and enhance the discrimination of visual features, facilitating seamless interaction between visual and textual branches. We believe this work provides valuable insights and potential directions for future research into the role of prompt learning in advancing CC-ReID.
While learnable prompts enable the model to autonomously construct and learn, have indeed demonstrated superior outcomes and a significant enhancement in performance compared to the manually crafted prompts used in CLIP, the prompts derived by the model may lack semantic clarity and remain illegible to human observers. As a result, it is not feasible to visually represent the learned content in a textual format. Thus, future research will need to address the interpretability of these prompts and investigate methodologies that enhance their comprehensibility while maintaining their efficacy.

### v2

@@
 Abstract
-Cloth-changing person re-identification (CC-ReID) aims to match individuals across multiple surveillance cameras despite variations in clothing.
-Existing methods typically focus on mitigating the effects of clothing changes or enhancing ID-relevant features but often struggle to capture complex semantic information.
-In this paper, we propose a novel prompt learning framework, Semantic Contextual Integration (SCI), for CC-ReID, which leverages the visual-text representation capabilities of CLIP to minimize the impact of clothing changes and enhance ID-relevant features.
-Specifically, we introduce Semantic Separation Enhancement (SSE) module, which uses dual learnable text tokens to separately capture confounding and clothing-related semantic information, effectively isolating ID-relevant features from distracting clothing semantics.
-Additionally, we develop a Semantic-Guided Interaction Module (SIM) that uses orthogonalized text features to guide visual representations, sharpening the model’s focus on distinctive ID characteristics.
-This integration enhances the model’s discriminative power and enriches the visual context with high-dimensional semantic insights.
+Cloth-changing person re-identification (CC-ReID) aims to match individuals across surveillance cameras despite variations in clothing.
+Existing methods typically mitigate the impact of clothing changes or enhance identity (ID)-relevant features, but they often struggle to capture complex semantic information.
+In this paper, we propose a novel prompt learning framework Semantic Contextual Integration (SCI), which leverages the visual-textual representation capabilities of CLIP to reduce clothing-induced discrepancies and strengthen ID cues.
+Specifically, we introduce the Semantic Separation Enhancement (SSE) module, which employs dual learnable text tokens to disentangle clothing-related semantics from confounding factors, thereby isolating ID-relevant features.
+Furthermore, we develop a Semantic-Guided Interaction Module (SIM) that uses orthogonalized text features to guide visual representations, sharpening the focus of the model on distinctive ID characteristics.
+This semantic integration improves the discriminative power of the model and enriches the visual context with high-dimensional insights.
 Extensive experiments on three CC-ReID datasets demonstrate that our method outperforms state-of-the-art techniques.
-The code will be released at github.
+The code will be released at https://github.com/hxy-499/CCREID-SCI.
 Index Terms:
 Person Re-Identification, Clothing Changes, Vision-Language Models, Prompt Learning, Semantic Integration.I Introduction
-Person re-identification (Re-ID) is a critical task that involves identifying individuals across different locations over time, with significant applications in video surveillance and smart city infrastructure [1, 2, 3].
-Traditional Re-ID methods [4, 5, 6, 7] focus on clothing features, such as texture and color, to improve performance.
-However, these methods become ineffective when individuals change clothes, highlighting the need for robust approaches that can handle clothing variations in real-world settings.
-To address the challenge of clothing changes in Re-ID (CC-ReID), extensive research has yielded promising results [8, 9].
-Existing methods can be broadly categorized into two categories: those aiming to reduce the influence of clothing and those focusing on enhancing ID-relevant information.
-The first category seeks to minimize the impact of clothing features, as illustrated in Fig.
-1(a).
-For instance, CAL [10] uses adversarial learning to reduce reliance on clothing by penalizing predictions based on the RGB modality, while AIM [11] employs causal intervention with a dual-branch model to mitigate clothing bias.
-The second category emphasizes ID-relevant features using auxiliary cues such as pose, gait, or human parsing networks.
-For example, GI-ReID [12] uses gait information to learn clothing-agnostic representations, and SCNet [13] leverages human body parsing network to learn identity-relevant features through semantic consistency constraints.
-Although these methods address clothing changes, they typically focus either on minimizing clothing influence or enhancing ID cues through explicit features like body contours.
-We argue that implicit positive factors such as hair, glasses, or backpacks are embedded in the visual features but are challenging to define explicitly or separately.
-As shown in Fig.
+Person re-identification (ReID) is the task of matching individuals across non-overlapping cameras over time, with important applications in video surveillance and smart city systems [1, 2, 3, 4].
+Traditional ReID methods [5, 6] exploit the appearance of clothing, e.g., texture and color, to distinguish identities (IDs), but their performance degrades when subjects change clothing.
+This limitation motivates the development of cloth-changing ReID (CC-ReID) techniques that remain robust under clothing variations.
+Existing CC-ReID approaches can be grouped into two categories: those that suppress clothing cues and those that enhance ID-relevant features (see Fig.
+1(a)).
+The first category seeks to minimize the influence of clothing: CAL [7] employs adversarial learning to penalize clothing-based predictions in the RGB modality, and AIM [8] uses causal intervention through a dual-branch model to mitigate clothing bias.
+The second category injects auxiliary ID cues, such as pose, gait, or human parsing, to strengthen discriminative features.
+For example, GI-ReID [9] leverages gait information to learn clothing-agnostic representations, while SCNet [10] imposes semantic consistency constraints using a human parsing network.
+Although these methods partially address clothing changes, they typically focus on minimizing negative clothing effects or relying on explicit cues such as body contours.
+We argue that visual representations inherently contain both negative factors (e.g., clothing styles that frequently change and can confuse the model) and positive factors (e.g., stable attributes such as hair, glasses, or backpacks that aid ID discrimination).
+As illustrated in Fig.
 1(b), our goal is to remove negative influences while making implicit positive factors more explicit.
-Recent advancements in visual-language learning have shown great potential in enhancing visual recognition and understanding.
-Contrastive Language-Image Pre-training (CLIP) [14], a prominent cross-modal pre-training model, has effectively bridged the gap between visual data and natural language, enabling more context-aware and interpretatively rich analysis.
-CLIP has achieved remarkable success in various downstream visual tasks [15, 16].
-In Re-ID, recent work [17, 18] demonstrates that linking visual content with corresponding language descriptions allows models to capture complex semantics associated with individuals.
-However, these studies have primarily focused on traditional Re-ID, applying CLIP to CC-ReID has not yet been explored.
-This motivates us to leverage prompt learning to reveal high-dimensional semantic information, thereby enhancing visual-textual connections and making implicit positive factors more explicit.
-In this paper, we propose the Semantic Contextual Integration (SCI) framework for CC-ReID, which utilizes CLIP to capture high-dimensional semantic features.
-Since CLIP struggles to selectively ignore aspects like clothing through negative prompts, we introduce a Semantic Separation Enhancement (SSE) module to minimize the influence of clothing by isolating semantic-level negative factors and preserving key positive features for CC-ReID.
-Specifically, we use dual learnable text tokens to separately capture confounding and clothing-related semantic information.
-We then apply an orthogonalization process on two distinct sets of text features to filter out clothing-related features from the confounding ones.
-SSE not only reduces the negative impact of clothing but also retains important positive semantic factors.
-These refined text features are used to guide the visual encoder.
-To further emphasize implicit positive factors in visual images, we introduce a Semantic-Guided Interaction Module (SIM), which directs the visual encoder’s extraction process using the filtered text features.
-This allows the model to focus on ID-relevant elements beyond explicit features such as body contours.
-By leveraging CLIP’s text-visual integration capabilities, our method enhances the interaction between the two modalities, improving visual feature refinement and multi-modal alignment.
-Our contributions are summarized fourfold:
+Recent advances in vision-language learning, notably Contrastive Language-Image Pre-training (CLIP) [11], have demonstrated strong abilities to bridge visual data and natural language, yielding context-aware representations beneficial for many downstream tasks [12, 13].
+In ReID, methods like CLIP-ReID [14] and Instruct-ReID [15] show that aligning images with descriptive language captures rich semantics beyond appearance alone.
+Inspired by these successes, we leverage prompt learning to extract high-dimensional semantic cues from CLIP, thereby amplifying positive factors without additional data preprocessing.
+In this paper, we introduce the Semantic Contextual Integration (SCI) framework for CC-ReID.
+SCI uses CLIP to capture rich semantic features and includes a Semantic Separation Enhancement (SSE) module isolates clothing-related negative factors at the text level while preserving key positive semantics for ReID.
+Specifically, SSE employs dual learnable text tokens to disentangle confounding semantics (both positive and negative) from clothing semantics, then orthogonalizes the token embeddings to filter out clothing factors.
+The refined text features guide visual encoding via our Semantic-Guided Interaction Module (SIM), focusing the model’s attention on ID-relevant cues beyond explicit contours.
+By integrating CLIP’s multimodal strengths, SCI enriches visual representations with high-dimensional semantic context and achieves superior performance on three standard CC-ReID benchmarks.
+Our contributions are summarized threefold:
 We propose the Semantic Contextual Integration (SCI) framework to leverage semantic information in CC-ReID, removing negative factors, emphasizing implicit positive elements, and refining visual representations.
@@
 We design the Semantic-Guided Interaction Module (SIM) to guide visual representations using refined text features, enhancing multi-modal integration and alignment.
-Extensive experiments on three standard CC-ReID datasets demonstrate the effectiveness of our approach, providing a robust solution to real-world Re-ID challenges.
 II Related Work
-II-A Cloth-Consistent Person Re-ID
-Person Re-ID under consistent clothing conditions has been extensively studied [19, 20, 21].
-These methods primarily rely on clothing color and texture to extract discriminative features.
-Feature representation learning in this context can be categorized into three types: global features, local features, and auxiliary information-based features.
-Global feature representation methods extract a single feature vector for each person image [22].
-Local feature representation methods aggregate part-based features to address misalignments, using techniques like human parsing [23] for semantic part detection, horizontal division [24] for structured feature aggregation, hard region-level and soft pixel-level attention [25] for optimized features, and IANet [26] for enhanced small visual cues.
-Auxiliary information-based methods incorporate additional data, such as semantic attributes [19] or generated/augmented samples [20], to enhance generalization by embedding richer contextual information.
-And ISP [27] locates human body parts and personal belongings at pixel-level for aligned with labels.
-Although effective in handling variations in appearance, background, and camera perspectives, these methods assume consistent clothing, leading to performance declines in scenarios with clothing changes, which are common in real-world applications.
-This limitation has spurred increased focus on addressing cloth-changing person Re-ID challenges.
-II-B Cloth-Changing Person Re-ID
-To advance CC-ReID, several datasets such as PRCC [28], LTCC [29], Celeb-reID [30], and VC-Clothes [31] have been introduced, providing diverse data for effective model evaluation.
-The core challenge in CC-ReID is to learn features that reliably identify individuals despite clothing variations.
-There are three primary approaches to this challenge.
-The first leverages auxiliary soft-biometric information.
-For example, CAMC [32] incorporates body shape semantics into ID-related features, enhancing robustness to clothing changes, while FSAM [33] learns fine-grained body shape features to complement clothing-independent features.
-M2Net [34] utilizes contour and human parsing information to learn features that are robust to appearance changes.
-The second approach involves disentangling features to isolate clothing-irrelevant information and reduce clothing bias.
-For instance, CAL [10] uses clothes-based adversarial loss to emphasize clothing-irrelevant features, and AIM [11] applies causal intervention to automatically eliminate clothing effects.
-The third approach employs data augmentation or adaptation techniques.
-Pos-Neg [8] and CCFA [9] enhance clothing color and texture diversity without additional data collection, improving robustness to clothing variations.
-RCSANet [35] explicitly constructs a clothing status, which is used to enhance the robustness of ID features for handling both cloth-changing and no-cloth-changing cases.
-Despite these advancements, existing methods primarily rely on visual inputs for model training, limiting their capacity for deep semantic understanding.
-In this paper, we propose a novel method that integrates visual and natural language information, utilizing CLIP to enhance the interaction of the visual content with corresponding textual descriptions.
+II-A Cloth-Consistent Person ReID
+Person ReID under consistent clothing conditions has been extensively studied [16].
+These methods exploit clothing appearance, such as color and texture, to extract discriminative features.
+Feature representation learning in this context can be divided into three categories: global features, local features, and auxiliary information-based features.
+Global methods extract a single feature vector per image [17], while local methods aggregate part-based representations to address misalignment (e.g., human parsing [18] or horizontal partitioning [19]).
+Auxiliary-information approaches incorporate extra cues, such as semantic attributes [20] or synthesized samples [21], to enrich context.
+Strong baselines in this domain include AGW [22] and TransReID [23].
+ISP [24] further refines alignment by locating body parts and carried items at the pixel level.
+Moreover, since person ReID is inherently formulated as a ranking problem, class imbalance can also have a significant impact on performance.
+Therefore, some works propose loss function modifications to handle data imbalance during training.
+For example, DRSL [25] optimizes the ReID model by enforcing positives to rank ahead of negatives based on distance, and further sorting positives by similarity score.
+Parameterized RV Loss [26] jointly optimizes retrieval and verification tasks by aligning loss functions with evaluation metrics, enabling automatic loss function search.
+Although these techniques handle variations in pose, background, and viewpoint, they assume identical clothing across views.
+Consequently, their performance degrades when subjects change clothing, motivating research on CC-ReID.
+II-B Cloth-Changing Person ReID
+Several CC-ReID datasets have been proposed, PRCC [27], LTCC [28], COCAS+ [29], and VC-Clothes [30], to evaluate models under clothing variations.
+The primary challenge is to learn features that remain reliable despite clothing changes.
+Existing CC-ReID methods follow three main strategies:
+II-B1 Auxiliary Soft-Biometric Cues
+CAMC [31] integrates body-shape semantics into ID features, and FSAM [32] extracts fine-grained shape information to complement clothing-independent cues.
+M2Net [33] uses contour and parsing maps for appearance-robust features, PGAL [34] aligns keypoints via pose estimation, FLAG [35] is proposed to explicitly extract appearance and gait information, and can be integrated with most existing video-based ReID methods, and CVSL [36] jointly learns body-shape embeddings and appearance features.
+II-B2 Feature Disentanglement
+CAL [7] employs adversarial loss to suppress clothing-related features, and AIM [8] uses causal intervention to remove clothing bias.
+DCR-ReID [37] disentangles and reconstructs feature components, LDF [38] separates ID, clothing, and unrelated factors via GANs, 3DInvarReID [39] disentangles features and reconstructs two-layer 3D body shapes, FIRe2 [40] augments images with fine-grained attributes, and MAL-F [41] learns invariant features from RGB, grayscale, and contour inputs using a ResTNet backbone.
+II-B3 Data Augmentation and Adaptation
+Pos-Neg [42] and CCFA [43] augment clothing color and texture diversity; RCSANet [44] constructs explicit clothing-status embeddings to bolster feature robustness; MCSC [45] applies meta-learning to address clothing-distribution shifts; Zhao et al.
+[46] model clothing changes as fine-grained domain shifts via graph relations; and DCLR [47] synthesizes multi-clothing images via diffusion and merges them into training data.
+Despite these advances, most CC-ReID methods rely solely on visual inputs, limiting their semantic depth.
+In this work, we integrate visual and textual modalities via CLIP to enrich semantic understanding and improve cloth-changing ReID.
 II-C Vision-Language Learning
-Vision-language pre-training (VLP) has recently shown significant promise in enhancing the performance of various downstream vision tasks by training models to associate images with textual information.
-CLIP [14], a leading VLP model, employs paired image and text encoders to bridge the gap between natural language and visual content.
-By using similarity learning, CLIP aligns textual and visual representations in a shared space, enhancing tasks such as image captioning and classification.
-Building on CLIP, recent advances in prompt learning, such as CoOp [36], transform prompt context words into learnable vectors.
-CoCoOp [37] extends this by generating input-conditional tokens for each image through a lightweight network.
-DenseCLIP [38] adapts this approach for dense prediction models, introducing pixel-text matching to utilize contextual information from a pre-trained language model.
-In person Re-ID, CLIP-ReID [17] is the first to use ID-specific learnable tokens and a two-stage training strategy to fully utilize text for specifying visual concepts.
-CCLNet [18] advances this approach with a novel prompt learning paradigm for unsupervised visible-infrared Re-ID.
-However, these studies primarily focus on scenarios with consistent clothing, leaving CLIP’s potential in CC-ReID largely unexplored.
-In this paper, we propose the CLIP-CCReID framework to mitigate the impact of clothing changes in CC-ReID.
+Vision-language pre-training (VLP) trains models to align images with text, improving downstream visual tasks.
+CLIP [11] uses paired image and text encoders to learn a shared embedding space via contrastive learning, benefiting applications such as captioning and classification.
+Prompt learning extends CLIP by making context tokens learnable.
+CoOp [48] transforms prompt words into trainable vectors, CoCoOp [49] generates input-conditional tokens per image, and DenseCLIP [50] applies pixel-text matching for dense prediction.
+In person ReID, CLIP-ReID [14] introduces ID-specific tokens and a two-stage training scheme.
+CCLNet [51] adapts prompt learning for unsupervised visible-infrared ReID.
+CSDN [52] employs bimodal descriptions to align visible and infrared features.
+RGANet [53] uses CLIP to locate informative body parts for occluded ReID.
+VGSG [54] groups text features semantically to address fine-grained misalignment.
+CCAFL [55] integrates CLIP-generated textual information into a novel semi-supervised framework to actively learn camera-agnostic pedestrian features.
+MIPL [56] guides CC-ReID with multiple common prompts, and CCAF [57] leverages CLIP’s knowledge to learn fine-grained, clothing-independent semantics.
+Although CLIP is not specifically pre-trained for ReID, its large-scale image-text training captures high-level semantics (e.g., body shape and context) remain invariant to clothing.
+We build on this by proposing CLIP-CCReID, which uses prompt learning to mitigate clothing changes without auxiliary models.
 III Proposed Method
 III-A Preliminaries and Overview
-The large-scale pre-trained vision-language model CLIP [14] demonstrates significant advantages in learning joint representations of images and text.
-CLIP consists of two encoders: a visual encoder and a text encoder .
-CLIP aligns the embedding spaces of visual and textual data during pre-training using a contrastive objective.
-For downstream classification tasks, a simple and effective method [36] constructs text prompts using a template like “a photo of a [CLS].”, where [CLS] is replaced with specific class names.
-Given an image, CLIP computes similarities between the image embedding and the text prompt embeddings, the process involves transforming the image and text into a shared high-dimensional space where their relationships can be quantitatively assessed.
-CLIP then selects the class corresponding to the highest similarity as the prediction.
-Building upon CLIP, CoOp [36] introduces learnable textual contexts, represented as “ [CLS].”, where () are learnable tokens.
-This approach enhances the model’s transferability and performance.
-Leveraging the capabilities of CLIP and CoOp, we employ these textual descriptions to achieve deeper semantic understanding for CC-ReID, as illustrated in Fig.
-2, enabling more robust and accurate modeling despite clothing changes.
+The large-scale vision-language model CLIP [11] learns joint image-text representations via a contrastive objective.
+It comprises a visual encoder and a text encoder , whose embedding spaces are aligned during pre-training.
+For downstream classification, CoOp [48] refines CLIP’s zero-shot prompts by introducing learnable context tokens.
+Specifically, given a template such as []1 []2 []M [CLS] where each []i () is a trainable vector and [CLS] is a class token, CLIP computes image-text similarities in the shared embedding space and predicts the class with maximal similarity.
+Building on these ideas, we leverage CLIP and CoOp to extract rich semantic prompts for CC-ReID (see Fig. 2), enabling robust ID modeling under clothing changes.
 III-B Semantic Separation Enhancement
-Since CLIP cannot learn negative semantics or ignore specific elements like clothing through negative prompts, we propose the Semantic Separation Enhancement (SSE) module to address this limitation.
-This module mitigates the impact of clothing variations by subtracting negative semantic elements, preserving only the high-dimensional positive features crucial for CC-ReID.
-The complete algorithm is summarized in Algorithm 1.
-We define dual textual prompts as follows:
+CLIP cannot explicitly ignore unwanted semantics (e.g., clothing) via negative prompts.
+To address this, we propose the Semantic Separation Enhancement (SSE) module, which isolates and removes clothing-related components from the high-dimensional text features, preserving only ID-relevant semantics.
+The overall procedure is detailed in Algorithm 1.
+We define two learnable prompts:
 where () are learnable tokens initialized randomly, and denotes the number of tokens.
@@
 The computation process is as follows:
-We then perform orthogonalization by subtracting the projection from , removing the component in the direction of at the textual level.
+We then perform orthogonalization by subtracting the projection of the clothing feature onto the .
+This operation removes the clothing component that is aligned with the ID direction.
 This process ensures that the features are aligned with the positive features while removing the influence of negative factors.
 The orthogonalization process is:
-where is the feature orthogonalized to align with and orthogonal to .
+where is after removing the part that is related to clothing, ensuring the feature focuses on positive information while reducing interference from clothing.
 To optimize the text features, we design a loss function that maximizes the similarity between and , while minimizing the similarity with .
@@
 III-C Semantic-Guide Interaction Module
-Previous work [10, 17, 39] shows that reducing clothing bias in visual or textual features improves performance in cloth-changing scenarios.
+Previous work [7, 14] shows that reducing clothing bias in visual or textual features improves performance in cloth-changing scenarios.
 However, these methods typically treat visual and textual branches independently, lacking interaction.
@@
 As illustrated in Fig.
-3, we employ a cross-attention mechanism in a Transformer decoder [40] to model interactions between visual and textual data.
-Instead of traditional self-attention, we use a non-local operation [41] to capture global context, ensuring the model can integrate information from distant but relevant positive parts of the visual image, which is crucial in cloth-changing scenarios.
+3, we employ a Transformer decoder [58] to model interactions between visual and textual data.
+To better capture long-range dependencies within the visual domain, we replace the conventional self-attention mechanism with a non-local operation [59], which allows the model to aggregate information from spatially distant but semantically relevant regions.
+This capability is particularly important for handling variations in clothing.
 During this stage, we update the visual encoder while keeping other components frozen.
@@
 where contains learnable parameters. is the final output feature map obtained by adding this context-aware feature to the original feature map.
-We further enhance interaction by using text embeddings as queries and visual embeddings as keys and values.
-This cross-attention mechanism refines visual features using text information, ensuring alignment with relevant semantic context:
-where is the dimensionality of the keys. This implementation strategically encourages the text features to find the most relevant visual clues.
+To refine visual representations with semantic guidance, we leverage textual embeddings to generate attention weights over visual features.
+This allows the model to selectively emphasize semantically relevant regions in the visual space, leading to more discriminative and context-aware visual representations:
+where is the dimensionality of the keys, is used to normalize the input features to stabilize training, and denotes a two-layer feed-forward network.
+This integrates visual features under the guidance of textual embeddings, and enables the model to extract the most informative visual cues aligned with the semantic content of the text.
 Finally, we update the visual features.
@@
 where is a learnable parameter controlling the degree of refinement applied to the visual features.
-This module employs an interaction-to-alignment approach, ensuring that text embeddings guide the visual encoding process, transforming implicit information into explicit visual cues.
-By emphasizing positive factors and reducing negative ones, this method enhances the model’s ability to consistently identify individuals across various clothing scenarios.
+This module adopts an interaction-to-alignment approach, where text embeddings guide the visual encoding process.
+By emphasizing positive cues and mitigating negative ones, this method improves the model’s ability to consistently recognize individuals across different clothing scenarios.
 For the objective function, we incorporate cross-entropy loss and clothes-based adversarial loss [7] to optimize the visual encoder:
 where is the true label for the -th sample, and is the predicted probability of the true label . The clothes-based adversarial loss is expressed as:
-where is the number of clothing categories, is the clothes classifier, is the weight for the -th class, is the feature of sample , and is a temperature parameter.
+where is the number of clothing categories, is the clothes classifier, is the weight for the -th class, is the feature of sample , is a temperature parameter, and is the set of clothes classes with the same ID.
+The objective of is to encourage the backbone to extract clothes-irrelevant features by penalizing its predictive power regarding clothing.
 To fully leverage the capabilities of CLIP, we calculate the image-to-text cross-entropy loss as:
 where label smoothing is applied to . The total loss for the SIM module is:
+By jointly optimizing , , and , the model captures ID features from easy samples (same clothes) and progressively learns to distinguish hard cases (same ID, different clothes) while aligning visual and textual representations.
+The weights of loss terms are basic terms [7, 14] and set as 1.
 IV Experimental Results
 IV-A Datasets and Evaluation Metrics
-We evaluate our proposed method on three standard cloth-changing datasets: LTCC [29], PRCC [28], and VC-Clothes [31], following CAL [10] and AIM [11].
-Additionally, we also evaluate our method on two traditional Re-ID datasets: Market-1501 [42], and MSMT17 [43], to validate its applicability.
-LTCC dataset contains 17,138 images of 152 IDs captured by 12 cameras.
-The training set comprises 9,576 images of 77 IDs, and the testing set includes 7,543 images (493 queries and 7,050 gallery images) of 75 IDs.
-This long-term Re-ID dataset features frequent clothing changes and varied environmental conditions, with each ID having between 2 and 14 outfits.
-PRCC dataset consists of 33,698 images of 221 IDs captured by three cameras.
-The training set contains 22,898 images of 150 IDs, while the testing set includes 10,800 images of 71 IDs.
-Each person has two outfits, with cameras A and B capturing the same clothes and camera C capturing a different outfit.
-VC-Clothes dataset is a synthetic dataset generated using the GTA5 game engine.
-It comprises 9,449 images of 256 IDs across both training and testing sets, featuring 1,241 clothing items across four camera views.
-Market-1501 dataset includes 1,501 individuals recorded by six different cameras.
-It consists of 12,936 images of 751 IDs for training and 19,732 images of 750 IDs for testing.
-The evaluation is conducted using the single query setting.
-MSMT17 dataset includes 1,041 individuals captured by 15 different cameras.
-The training set contains 32,621 images of 1,041 IDs, while the testing set includes 51,027 images of 11,659 IDs.
-This dataset is notably large and complex.
-We use the cumulative matching characteristic (CMC) curve and mean average precision (mAP) as evaluation metrics.
-Rank- in CMC measures the likelihood of retrieving the correct ID in the top- results, while mAP captures the average retrieval performance across all queries.
-Evaluation is conducted under three settings: 1) General setting: Excludes same-ID and same-camera samples, using both cloth-changing and clothing-consistent samples.
-2) Same-clothes setting: Excludes same-ID and same-camera samples, using only clothing-consistent samples.
-3) Cloth-changing setting: Excludes same-ID, same-camera, and same-clothing samples, using only cloth-changing samples.
+We evaluate our method on three standard cloth-changing datasets: LTCC [28], PRCC [27], and VC-Clothes [30], following the protocols of CAL [7] and AIM [8].
+Additionally, to validate generalization, we test on two conventional ReID datasets: Market1501 [60] and MSMT17 [61].
+LTCC contains 17,138 images of 152 IDs captured by 12 cameras.
+The training set comprises 9,576 images of 77 IDs; the testing set includes 7,543 images (493 queries and 7,050 gallery) of 75 IDs.
+This long-term dataset features frequent clothing changes and varied environmental conditions, with each ID wearing between two and fourteen outfits.
+PRCC consists of 33,698 images of 221 IDs from three cameras.
+Its training set has 22,898 images of 150 IDs, and the testing set has 10,800 images of 71 IDs.
+Each person appears in two outfits: cameras A and B share the same clothing, while camera C uses a different outfit.
+VC-Clothes is a synthetic dataset generated via the GTA5 engine.
+It contains 9,449 images of 256 IDs across four camera views, with 1,241 distinct clothing items distributed between training and testing sets.
+Market1501 includes 12,936 training images of 751 IDs and 19,732 testing images of 750 IDs, captured by six cameras under the single-query evaluation protocol.
+MSMT17 comprises 32,621 training images of 1,041 IDs and 51,027 testing images of 11,659 IDs across 15 cameras, it is one of the largest and most challenging ReID benchmarks.
+We employ the cumulative matching characteristic (CMC) curve and mean average precision (mAP) as evaluation metrics.
+Rank- in CMC measures the probability of the correct match appearing among the top- retrieved images, while mAP reflects the overall retrieval performance across all queries.
+Experiments are conducted under three protocols: 1) General setting: exclude same-ID and same-camera samples, using both cloth-changing and clothing-consistent pairs; 2) Same-clothes setting: exclude same-ID and same-camera samples, using only clothing-consistent pairs; 3) Cloth-changing setting: exclude same-ID, same-camera and same-clothing samples, using only cloth-changing pairs.
 IV-B Implementation Details
-We adopt a modified ResNet-50 pre-trained on CLIP as the backbone for feature extraction.
-Images are resized to pixels, with a batch size of 64.
-Training employs the Adam optimizer [44] with random horizontal flipping, cropping, and erasing [45] for data augmentation.
-A global attention pooling layer reduces the feature dimension from 2048 to 1024, matching the text feature dimension, which is scaled from 512 to 1024.
-In the first stage, two prompt learners are trained for 60 epochs on LTCC, VC-Clothes, Market-1501, and MSMT17, and 20 epochs on PRCC, with an initial learning rate of adjusted using a cosine decay schedule.
-In the second stage, the visual encoder is trained for 120 epochs on LTCC, VC-Clothes, Market-1501, and MSMT17, and 20 epochs on PRCC, with learning rates initially set to and reduced by a factor of 10 at the 40th and 70th epochs.
-All experiments are conducted on the PyTorch platform using a single A100 GPU.
+We adopt two backbones for feature extraction: a modified ResNet-50 and ViT-B/16 pre-trained on CLIP.
+All images are resized to pixels, with a batch size of 64.
+We use the Adam optimizer [62] and apply random horizontal flipping, cropping, and erasing [63] for data augmentation.
+For the CNN branch, a global attention pooling layer to reduce the feature dimension from 2048 to 1024, matching the text feature dimension (scaled from 512 to 1024).
+For the transformer branch, we employ 12 transformer layers with hidden size 768; a linear projection reduces the image feature dimension from 768 to 512 to align with the text features.
+Training proceeds in two stages: 1) Stage 1 (prompt learning): train two prompt learners for 60 epochs (ResNet-50) or 120 epochs (ViT-B/16) with an initial learning rate of and cosine decay.
+2) Stage 2 (visual encoder fine-tuning): train the visual encoder for 120 epochs (ResNet-50) or 30 epochs (ViT-B/16), starting at and reducing by a factor of 10 at epochs 40 and 70.
+All experiments are implemented in PyTorch and run on a single NVIDIA A100 GPU.
 IV-C Comparison with State-of-the-Art Methods
-We compare our method with state-of-the-art techniques on LTCC, PRCC, and VC-Clothes, including traditional Re-ID methods HACNN [25], PCB [24], IANet [26], and ISP [27], and specialized CC-ReID methods FSAM [33], RCSANet [35], CAL [10], GI-ReID [12], M2Net [34], and AIM [11].
+We compare SCI against state-of-the-art methods on LTCC, PRCC and VC-Clothes, including general ReID approaches (ISP [24], CLIP-ReID [14], and Instruct-ReID [15]) and specialized cloth-changing methods (FSAM [32], RCSANet [44], CAL [7], GI-ReID [9], M2Net [33], AIM [8], Pos-Neg [42], ACID [64], 3DInvarReID [39], DCR-ReID [37], IMS+GEP [46], AFL [4], CVSL [36], MCSC [45], MAL-F [41], and DLCR [47]).
 Methods | Venue | LTCC | PRCC | VC-Clothes | |||||||||
@@
 R-1 | mAP | R-1 | mAP | R-1 | mAP | R-1 | mAP | R-1 | mAP | R-1 | mAP | ||
-| HACNN [25] | CVPR ’18 | 60.2 | 26.7 | 21.6 | 9.3 | 82.5 | 84.8 | 21.8 | 23.2 | 68.6 | 69.7 | 49.6 | 50.1 |
-| PCB [24] | ECCV ’18 | 65.1 | 30.6 | 23.5 | 10.0 | 99.8 | 97.0 | 41.8 | 38.7 | 87.7 | 74.6 | 62.0 | 62.2 |
-| IANet [26] | CVPR ’19 | 63.7 | 31.0 | 25.0 | 12.6 | 99.4 | 98.3 | 46.3 | 45.9 | - | - | - | - |
 ISP [24] | ECCV’20 | 66.3 | 29.6 | 27.8 | 11.9 | 92.8 | - | 36.6 | - | 94.5 | 94.7 | 72.0 | 72.1 |
@@
 RCSANet [44] | ICCV’21 | - | - | - | - | 100.0 | 97.2 | 50.2 | 48.6 | - | - | - | - |
+GI-ReID [9] | CVPR’22 | 63.2 | 29.4 | 23.7 | 10.4 | - | - | - | - | - | - | 64.5 | 57.8 |
+Pos-Neg [42] | TIP’22 | 75.7 | 37.0 | 36.2 | 14.4 | - | - | 54.9 | 65.8 | - | - | - | - |
 CAL [7] | CVPR’22 | 73.4 | 39.4 | 38.0 | 17.2 | 100.0 | 99.6 | 54.7 | 55.4 | 93.1 | 88.3 | 82.6 | 81.7 |
-| GI-ReID [12] | CVPR’22 | 63.2 | 29.4 | 23.7 | 10.4 | - | - | - | - | - | - | 64.5 | 57.8 |
 M2Net [33] | ACM MM’22 | - | - | - | - | 99.5 | 99.1 | 59.3 | 57.7 | - | - | - | - |
+ACID [64] | TIP’23 | 65.1 | 30.6 | 29.1 | 14.5 | 99.1 | 99.0 | 55.4 | 66.1 | 95.1 | 94.7 | 84.3 | 74.2 |
+CLIP-ReID [14] | AAAI’23 | 73.0 | 36.5 | 38.3 | 15.9 | 100.0 | 99.5 | 57.0 | 55.3 | 93.1 | 86.9 | 85.9 | 79.3 |
+3DInvarReID [39] | ICCV’23 | - | - | 40.9 | 18.9 | - | - | 56.5 | 57.2 | - | - | - | - |
+DCR-ReID [37] | TCSVT’23 | 76.1 | 42.3 | 41.1 | 20.4 | 100.0 | 99.7 | 57.2 | 57.4 | - | - | - | - |
 AIM [8] | CVPR’23 | 75.7 | 41.3 | 41.8 | 17.9 | 100.0 | 99.7 | 56.3 | 56.5 | 91.2 | 84.9 | 81.0 | 75.7 |
-| CLIP-ReID [17] | AAAI’23 | 73.0 | 36.5 | 38.3 | 15.9 | 100.0 | 99.5 | 57.0 | 55.3 | 93.1 | 86.9 | 85.9 | 79.3 |
-| SCI (Ours) | – | 75.7 | 40.6 | 42.1 | 18.6 | 99.6 | 97.7 | 59.8 | 56.2 | 94.9 | 89.2 | 89.2 | 83.1 |
-As shown in Table I, CC-ReID methods significantly outperform general Re-ID methods, primarily because general methods rely heavily on clothing for identification, which is less effective in long-term surveillance with clothing changes.
-Our SCI method achieves competitive results by effectively eliminating ID-irrelevant features and enhancing positive ID elements in visual images.
-On LTCC, SCI surpasses the baseline [17] in the cloth-changing setting, with improvements of 3.8% in Rank-1 accuracy and 2.7% in mAP.
-Although our mAP is slightly lower than AIM [11] in the general setting, this is likely due to LTCC’s inherent challenges, where models often rely on clothing information.
-By filtering out such information, SCI experiences minor declines in non-cloth-changing scenarios, but these are negligible given the overall benefits.
-Notably, SCI achieves these results without incorporating additional modalities or modules, unlike other methods.
-The primary advantage of our method lies in its ability to avoid carrying clothing-relevant information, which is often unavoidable in methods relying on auxiliary data such as poses and gaits.
-Compared with non-auxiliary-based methods, SCI outperforms CAL by 12.7% in Rank-1 and 6.3% in mAP in the cloth-changing setting on VC-Clothes.
-On PRCC, SCI also surpasses AIM by 3.5% in Rank-1 accuracy for clothing changes.
-Although the differences on other metrics are less pronounced due to PRCC’s simpler structure (only two outfits per ID), SCI still shows notable improvements.
-On VC-Clothes, SCI demonstrates significant gains in the cloth-changing setting, with a 7.4% higher mAP than AIM, underscoring the effectiveness of our approach in challenging scenarios.
+Instruct-ReID [15] | CVPR’24 | 75.8 | 52.0 | - | - | - | - | 54.2 | 52.3 | - | - | 89.7 | 78.9 |
+AFL [4] | TMM’24 | 74.4 | 39.1 | 42.1 | 18.4 | 100.0 | 99.7 | 57.4 | 56.5 | 93.9 | 88.3 | 82.5 | 83.0 |
+MCSC [45] | TIP’24 | 73.9 | 40.2 | 42.2 | 19.4 | 99.8 | 99.8 | 57.8 | 57.3 | 93.2 | 87.9 | 83.3 | 83.2 |
+IMS+GEP [46] | TMM’24 | - | - | 43.4 | 18.2 | 99.7 | 99.8 | 57.3 | 65.8 | 94.7 | 94.9 | 81.8 | 81.7 |
+CVSL [36] | WACV’24 | 76.4 | 41.9 | 44.5 | 21.3 | 97.5 | 99.1 | 57.5 | 56.9 | - | - | - | - |
+MAL-F [41] | TITS’25 | 37.5 | 16.2 | 38.8 | 17.0 | - | - | 57.1 | 57.7 | - | - | - | - |
+DLCR [47] | WACV’25 | - | - | 41.3 | 19.6 | - | - | 66.5 | 63.0 | - | - | 87.1 | 81.1 |
+SCI w/ ResNet-50 | Ours | 75.7 | 40.6 | 42.1 | 18.6 | 99.6 | 97.7 | 59.8 | 56.2 | 94.9 | 89.2 | 89.2 | 83.1 |
+SCI w/ ViT-B | Ours | 79.9 | 44.4 | 45.9 | 21.6 | 100.0 | 99.4 | 62.8 | 60.0 | 95.4 | 91.6 | 91.7 | 87.5 |
+As shown in Table I, SCI consistently outperforms baselines by effectively suppressing ID-irrelevant features and reinforcing discriminative ID cues.
+IV-C1 CNN-Based Model
+On LTCC (cloth-changing), SCI improves Rank-1 by 3.8% and mAP by 2.7% over CLIP-ReID [14].
+Although SCI’s mAP is marginally lower than AIM [8] in the general setting, this is expected: by filtering out clothing information, SCI may sacrifice some performance when clothing cues are reliable, but gains robustness under cloth variation.
+On PRCC, SCI surpasses AIM by 3.5% in Rank-1 accuracy under clothing changes. Due to PRCC’s simpler two-outfit per ID setup, improvements on other metrics are relatively smaller but still notable.
+On VC-Clothes, SCI achieves a substantial mAP gain of 7.4% over AIM in the cloth-changing setting, further validating its effectiveness in challenging scenarios.
+IV-C2 ViT-Based Model
+With ViT-B/16, SCI attains near state-of-the-art performance, especially on LTCC, reaching 45.9% Rank-1 under severe clothing variations.
+Crucially, these results are obtained without auxiliary models or additional preprocessing, distinguishing SCI from methods that rely on external components.
 IV-D Ablation Studies and Analysis
-We conduct ablation studies to assess the effectiveness of the components in SCI and analyze parameter impacts.
-IV-D1 Effectiveness of Semantic Separation Enhancement
+We perform ablation studies to assess each module’s contribution and analyze key parameters.
+Backbone | Baseline | SSE | SIM | LTCC | PRCC | VC-Clothes | |||||||||
+General | Cloth-changing | Same-clothes | Cloth-changing | General | Cloth-changing | ||||||||||
+R-1 | mAP | R-1 | mAP | R-1 | mAP | R-1 | mAP | R-1 | mAP | R-1 | mAP | ||||
+ResNet-50 | ● | ○ | ○ | 73.0 | 36.5 | 38.3 | 15.9 | 100.0 | 99.5 | 57.0 | 55.3 | 93.1 | 86.9 | 85.9 | 79.3 |
+● | ● | ○ | 76.5 | 39.7 | 39.8 | 18.3 | 100.0 | 98.4 | 59.0 | 54.7 | 94.1 | 88.0 | 87.6 | 81.1 | |
+● | ○ | ● | 72.8 | 36.3 | 37.8 | 16.1 | 99.7 | 98.0 | 52.4 | 49.9 | 94.0 | 88.9 | 88.4 | 82.4 | |
+● | ● | ● | 75.7 | 40.6 | 42.1 | 18.6 | 99.6 | 97.7 | 59.8 | 56.2 | 94.9 | 89.2 | 89.2 | 83.1 | |
+ViT-B | ● | ○ | ○ | 79.5 | 43.6 | 41.6 | 19.4 | 99.9 | 99.2 | 60.9 | 58.5 | 94.4 | 89.8 | 90.6 | 84.7 |
+● | ● | ○ | 76.3 | 43.9 | 43.4 | 21.0 | 100.0 | 99.3 | 62.3 | 58.7 | 94.6 | 89.8 | 91.4 | 84.8 | |
+● | ○ | ● | 77.5 | 43.9 | 42.9 | 20.5 | 99.8 | 99.2 | 61.4 | 59.1 | 94.3 | 90.4 | 90.9 | 85.7 | |
+● | ● | ● | 79.9 | 44.4 | 45.9 | 21.6 | 100.0 | 99.4 | 62.8 | 60.0 | 95.4 | 91.6 | 91.7 | 87.5 |
+IV-D1 Impact of Semantic Separation Enhancement
 We first evaluate the baseline without additional modules, where the model relies solely on a prompt “a photo of a person” for classification.
 This approach shows reduced performance.
-Introducing the SSE module, which uses two prompts (“a photo of a person” and “a photo of the clothes”) to learn features separately at the textual level, improves Rank-1 and mAP by around 2.0% across datasets, demonstrating its ability to filter clothing features.
+Introducing the SSE module, which uses two prompts (“a photo of a person” and “a photo of the clothes”) to learn features separately at the textual level.
+We observe consistent improvements of approximately 2.0% in both Rank-1 accuracy and mAP across datasets for both CNN-based and ViT-based models, as shown in Table II.
+This demonstrates the effectiveness of the SSE module in filtering clothing features.
 These results indicate: 1) The original generic textual representations are confounding, containing both ID and clothing features, leading the model to focus excessively on cloth-relevant aspects.
@@
 These findings demonstrate the effectiveness of setting specific prompts for each factor and the robustness of our SSE module against both cloth-changing and general scenarios.
-| Baseline | SSE | SIM | LTCC | PRCC | VC-Clothes | |||||||||
-| General | Cloth-changing | Same-clothes | Cloth-changing | General | Cloth-changing | |||||||||
-| R-1 | mAP | R-1 | mAP | R-1 | mAP | R-1 | mAP | R-1 | mAP | R-1 | mAP | |||
-| ● | ○ | ○ | 73.0 | 36.5 | 38.3 | 15.9 | 100.0 | 99.5 | 57.0 | 55.3 | 93.1 | 86.9 | 85.9 | 79.3 |
-| ● | ● | ○ | 76.5 | 39.7 | 39.8 | 18.3 | 100.0 | 98.4 | 59.0 | 54.7 | 94.1 | 88.0 | 87.6 | 81.1 |
-| ● | ○ | ● | 72.8 | 36.3 | 37.8 | 16.1 | 99.7 | 98.0 | 52.4 | 49.9 | 94.0 | 88.9 | 88.4 | 82.4 |
-| ● | ● | ● | 75.7 | 40.6 | 42.1 | 18.6 | 99.6 | 97.7 | 59.8 | 56.2 | 94.9 | 89.2 | 89.2 | 83.1 |
-IV-D2 Effectiveness of Semantic-Guided Interaction Module
+IV-D2 Impact of Semantic-Guided Interaction Module
 We further assess SIM by integrating it with the baseline model.
-As shown in Table II, using SIM alone often degrades performance, particularly in the cloth-changing scenario on PRCC, where Rank-1 drops by 4.6% and mAP decreases from 55.3% to 49.9%.
+As shown in Table II, using SIM alone often degrades performance based on CNN, particularly in the cloth-changing scenario on PRCC, where Rank-1 drops by 4.6% and mAP decreases from 55.3% to 49.9%.
 The decline in performance in this set of ablation experiments is primarily due to the absence of the SSE module.
@@
 (a) Baseline (stage1) | (b) Baseline (stage2) | (c) Ours (stage1) | (d) Ours (stage2) |
-IV-D3 Feature Distribution Analysis
+IV-D3 Feature Distribution via t-SNE
 To illustrate SCI’s effectiveness, we employ t-SNE visualization [65], as depicted in Fig.
@@
 This visualization plots the statistical distribution for 20 randomly selected categories from LTCC, comparing the latent space distributions at different stages of the baseline model with our approach.
-From Figs.
+From Fig.
 4(a) and (c), it is observable that both the baseline and our method exhibit disorganized feature distributions with blurred class boundaries.
 During this stage, only the prompts are trained while the visual and text encoders remain frozen, limiting their capability to effectively extract diverse representations.
-In contrast, as shown in Figs.
+In contrast, as shown in Fig.
 4(b) and (d), the feature distributions for both methods become clearer, indicating the effectiveness of the representations derived from the text-image similarity computations.
@@
 Therefore, this visualization not only serves as compelling validation of our capability to effectively extract features and discriminate IDs but also highlights our method’s potential to significantly advance the field in cloth-changing scenarios.
-IV-D4 Visualization of Feature Maps
-To illustrate the distinctions between the baseline [17] and SCI, we present heatmaps of regions of interest in Fig.
-5, demonstrating performance across LTCC, VC-Clothes, and PRCC.
+IV-D4 Qualitative Prompt-Guided Refinement
+To illustrate the distinctions between the baseline [14] and SCI, we present in Fig.
+5 the attention maps activated by different prompts, highlighting how prompts guide the refinement of visual features across LTCC, VC-Clothes, and PRCC.
+It demonstrates that the learned prompts are capable of guiding the network to focus on relevant semantic areas.
 Fig.
@@
 These visualizations provide insights into the unique differences in feature activation patterns.
+Moreover, while SCI leverages prompt learning to jointly model ID and clothing information, it fundamentally differs from attribute-based approaches, as it neither explicitly recognizes attributes nor relies on additional attribute detectors.
 As shown in Fig.
@@
 We are aware of this issue and will continue to optimize our model to address it.
-| (a) R-1 Accuracy on cloth-changing | (b) mAP Accuracy on cloth-changing | (c) R-1 Accuracy on general | (d) mAP Accuracy on general |
-IV-D5 Comparative Analysis of Mechanisms in SIM
+(a) distribution. | (b) distribution. | (c) distribution. |
+IV-D5 Visualization of Feature Similarity
+In Fig.
+7, we visualize the similarity distributions of features between the text prompts and the images.
+Fig.
+7(a) is derived from the prompt in Eq.
+1, while the distribution is relatively well-formed, intra-class similarity remains dispersed.
+Fig.
+7(b) shows the distribution from the prompt in Eq.
+2, which learns clothing-related semantics.
+However, due to significant clothing variations over time for the same individual, intra-class differences outweigh inter-class differences.
+Fig.
+7(c) shows the distribution after orthogonalizing the features, which achieves a clearer separation between intra- and inter-class similarity, indicating improved feature discrimination.
+This demonstrates the effectiveness of dual textual prompts and orthogonalization in enhancing feature alignment.
+(a) R-1 on cloth-changing | (b) mAP on cloth-changing | (c) R-1 on general | (d) mAP on general |
+IV-D6 Comparative Analysis of Mechanisms in SIM
 As shown in Fig.
-7, we evaluate the effectiveness of the non-local operation [41] compared to self-attention mechanisms [40] in leveraging textual information obtained from the SSE module to guide visual representations.
-In cloth-changing scenarios, the non-local mechanism yields superior performance, achieving 42.1% in Rank-1 accuracy and 18.6% in mAP, as depicted in Figs.
-7(a) and (b), thereby outperforming the self-attention mechanism by margins of 4.6% and 2.4%, respectively.
+8, we evaluate the effectiveness of the non-local operation [59] compared to self-attention mechanisms [58] in leveraging textual information obtained from the SSE module to guide visual representations.
+In cloth-changing scenarios, the non-local mechanism yields superior performance, achieving 42.1% in Rank-1 accuracy and 18.6% in mAP, as depicted in Fig.
+8(a) and (b), thereby outperforming the self-attention mechanism by margins of 4.6% and 2.4%, respectively.
 This underscores the non-local mechanism’s enhanced capability to capture comprehensive global context, which is vital for integrating information from disparate yet significant segments of the visual field in cloth-changing scenarios.
-In general scenarios, as illustrated in Figs.
-7(c) and (d), both mechanisms exhibit comparable efficacy, with the non-local mechanism slightly edging out self-attention by a 2.0% increase in mAP.
-This suggests the non-local mechanism is robust and can be effectively deployed for person Re-ID tasks.
-IV-D6 Loss Functions Analysis
-We conduct experimental evaluations to explore the effectiveness of (9), as demonstrated in Fig.
-8(a).
+In general scenarios, as illustrated in Fig.
+8(c) and (d), both mechanisms exhibit comparable efficacy, with the non-local mechanism slightly edging out self-attention by a 2.0% increase in mAP.
+This suggests the non-local mechanism is robust and can be effectively deployed for person ReID tasks.
+IV-D7 Parameter Sensitivity
+We conduct experimental evaluations to explore the effectiveness of Eq.
+8, as demonstrated in Fig.
+9(a).
 This loss function significantly improves performance by optimizing for higher ID similarity and reducing clothing similarity, making it particularly effective for applications that need to distinguish personal ID from external features like clothing, relevant in tasks such as cloth-changing scenarios.
 It helps the model focus on positive factors while ignoring negative variations such as clothing changes, thereby enhancing adaptability and accuracy in complex environments.
+IV-D8 Generalization to Cloth-Consistent Scenarios
+To validate the applicability of our method in cloth-consistent scenarios, we conducted experiments using our cloth-changing approach on two traditional ReID datasets: Market1501 and MSMT17.
+For ease of comparison, we also incorporated cloth-consistent methods: AGW [22], TransReID [23], CLIP-ReID [14], and Instruct-ReID [15].
+This enables a comprehensive evaluation of the performance in varying scenarios.
 Methods | Venue | Market1501 | MSMT17 | |||
@@
 Cloth- consistent | AGW [22] | TPAMI’21 | 95.1 | 87.8 | 68.3 | 49.3 |
-| reID-NAS [49] | TNNLS’22 | 95.1 | 85.7 | 79.5 | 53.3 | |
+TransReID [23] | ICCV’21 | 95.2 | 89.5 | 86.2 | 69.4 | |
 CLIP-ReID [14] | AAAI’23 | 95.7 | 89.8 | 84.4 | 63.0 | |
-| IRM [50] | CVPR’24 | 96.5 | 93.5 | 86.9 | 72.4 | |
+CMT [16] | TCSVT’24 | 95.8 | 88.9 | 78.5 | 51.3 | |
+Instruct-ReID [15] | CVPR’24 | 96.5 | 93.5 | 86.9 | 72.4 | |
 Cloth- changing | CAL [7] | CVPR’22 | 85.0 | 67.8 | 65.1 | 38.4 |
 AIM [8] | CVPR’23 | 70.3 | 48.0 | 26.6 | 12.0 | |
-| SCI (Ours) | – | 85.7 | 68.6 | 73.5 | 47.4 | |
-IV-D7 Cloth-consistent Re-ID Scenario
-To validate the applicability of our method in cloth-consistent scenarios, we conducted experiments using our cloth-changing approach on two traditional ReID datasets: Market1501 and MSMT17.
-For ease of comparison, we also incorporated cloth-consistent methods: AGW [48], reID-NAS [49], CLIP-ReID [17], and IRM [50].
-This enables a comprehensive evaluation of the performance in varying scenarios.
-As shown in Table III (cloth-consistent), existing Re-ID methods designed for traditional datasets have achieved progressively better performance, even on larger-scale datasets like MSMT17.
-While our method is specifically designed to address the more realistic scenario of clothing changes, we are also interested in evaluating its performance on conventional datasets.
+3DInvarReID [39] | ICCV’23 | 95.1 | 97.9 | 80.8 | 59.1 | |
+AFL [4] | TMM’24 | 95.5 | 88.8 | 80.9 | 58.1 | |
+FIRe2 [40] | TIFS’24 | 95.4 | 87.7 | 79.7 | 56.2 | |
+SCI w/ ResNet-50 | Ours | 94.4 | 86.5 | 81.4 | 59.5 | |
+SCI w/ ViT-B | Ours | 95.2 | 89.4 | 89.6 | 74.7 |
+Methods | Venue | Input Resolution | Params (M) | GFLOPs | Training Time (min) | FPS | R-1 | ||
+LTCC | PRCC | VC-Clothes | |||||||
+CAL [7] | CVPR’22 | 23.5 | 9.2 | 41 | 538 | 38.0 | 54.7 | 82.6 | |
+CLIP-ReID [14] | AAAI’23 | 77.6 | 10.6 | 67 | 314 | 38.3 | 57.0 | 85.9 | |
+3DInvarReID [39] | ICCV’23 | 217.6 | – | 241 | – | 40.9 | 56.5 | - | |
+AIM [8] | CVPR’23 | 72.7 | 9.2 | 52 | 754 | 41.8 | 56.3 | 81.0 | |
+SCI w/ ResNet-50 | Ours | 93.9 | 10.6 | 78 | 302 | 42.1 | 59.8 | 89.2 | |
+SCI w/ ViT-B | Ours | 143.7 | 26.7 | 46 | 80 | 45.9 | 62.8 | 91.7 |
+As shown in Table III (cloth-consistent), existing ReID methods designed for traditional datasets have achieved progressively better performance, even on larger-scale datasets like MSMT17.
+Although our proposed SCI framework is primarily designed to address the more practical scenario of clothing changes, we also evaluate its effectiveness on conventional datasets to assess its generalization ability.
 From Table III (cloth-changing), it can be observed that our SCI method achieves the best performance on both Market1501 and MSMT17 among cloth-changing methods.
-Although there is still a gap compared to cloth-consistent methods, this highlights an important direction for future research and improvement.
-IV-D8 Visualization of Retrieved Examples
+In particular, the ViT-based SCI achieves a Rank-1 accuracy of 89.6% on MSMT17, outperforming Instruct-ReID by 2.7% in Rank-1 and 2.3% in mAP.
+These results underscore the strong generalization capability of our approach.
+IV-D9 Visualization of Retrieved Examples
 We present the visualization of the top-10 retrieved results from randomly selected query examples on LTCC in Fig.
-9.
+10.
 It observes that the retrieval results are predominantly optimistic, reflecting a high level of accuracy and relevance in most cases.
-However, some incorrect results still persist, which could be due to challenges such as varying viewpoints, low lighting conditions, or occlusions within the images.
-These issues underscore potential areas for future research.
-Despite these challenges, the overall result strongly supports the effectiveness of our SCI, demonstrating its robustness across a variety of scenarios.
-IV-D9 Parameter Selection
-We evaluate the parameters and specified in (9).
+However, some incorrect results still persist, as shown in Fig.
+11, mainly due to limited views and highly similar clothing colors, which make retrieval harder.
+This problem is common when people wear uniforms or dark winter clothes.
+Further improvements are needed to enhance the model’s ability to distinguish individuals with similar appearances.
+IV-D10 Parameter Selection
+We evaluate the parameters and specified in Eq.
+8.
 The results for LTCC are displayed in Fig.
-8(b).
+9(b).
 We determine that and provide the best performance in most scenarios, optimizing the balance between enhancing ID and reducing clothing similarity, while feature orthogonalization minimizes interference between ID and clothing features.
@@
 This analysis underscores the effectiveness and adaptability of our proposed method in real-world surveillance applications.
-IV-D10 Efficiency Analysis
-To verify the efficiency and performance of SCI, we analyze several computational and training metrics across different methods, detailed in Table IV.
-AIM [11] utilizes a dual-branch ResNet model, achieving an R-1 accuracy of 41.8%, which surpasses CAL [10], another ResNet-based model.
-Although AIM requires higher GFLOPs compared to CAL, it significantly outperforms CAL in terms of accuracy with the same parameter count.
+IV-D11 Efficiency Comparison
+Table IV presents a comprehensive comparison of parameter size, computational complexity (GFLOPs), training time, inference speed (FPS), and Rank-1 accuracy across representative ReID methods.
+AIM [8] utilizes a dual-branch ResNet model, achieving an Rank-1 accuracy of 41.8%, which surpasses CAL [7], another CNN-based model.
+Although AIM requires higher parameters compared to CAL, it significantly outperforms CAL in terms of accuracy with the same parameter count.
 In CLIP-ReID [14] and our method, we integrate textual modality information, transitioning from a single-modality to a multimodal approach.
 This justifies the observed increases in both GFLOPs and parameters, which remain within acceptable limits.
-Our SCI achieves an R-1 accuracy of 42.1%, demonstrating the value of incorporating textual information.
-This allows the model to better distinguish between different IDs, especially in challenging scenarios.
-Overall, our method provides a good balance between computational cost and performance enhancements.
+Our SCI w/ ResNet-50 incorporates a prompt learning strategy and achieves 42.1% Rank-1 on LTCC, surpassing CLIP-ReID by 3.8%, while maintaining similar computational efficiency (10.6 GFLOPs vs.
+10.6, and 302 FPS vs.
+314).
+This highlights the effectiveness of our design in utilizing textual guidance without introducing significant overhead.
+Furthermore, our SCI w/ ViT-B achieves the best performance across all datasets, achieving 45.9% Rank-1 accuracy on LTCC, 62.8% on PRCC, and 91.7% on VC-Clothes.
+Although it has higher computational requirements due to the ViT backbone (143.7M parameters, 26.7 GFLOPs), the performance gains are substantial.
+Overall, SCI achieves a good trade-off between accuracy and efficiency, making it well-suited for practical deployment in real-world scenarios.
 V Conclusion and Limitation
-In this study, we propose the Semantic Contextual Integration (SCI) network, a novel method incorporates prompt learning to address cloth-changing person re-identification (CC-ReID).
-The SCI network includes a Semantic Separation Enhancement (SSE) module that generates ID-specific and clothing prompts to effectively isolate clothing semantics (negative factors) from ID information, thereby preserving high-dimensional semantic details (positive factors) embedded in visual images.
-Additionally, the Semantic-Guided Interaction Module (SIM) leverages textual cues to guide and enhance the discrimination of visual features, facilitating seamless interaction between visual and textual branches.
-We believe this work provides valuable insights and potential directions for future research into the role of prompt learning in advancing CC-ReID.
-While learnable prompts enable the model to autonomously construct and learn, have indeed demonstrated superior outcomes and a significant enhancement in performance compared to the manually crafted prompts used in CLIP, the prompts derived by the model may lack semantic clarity and remain illegible to human observers.
-As a result, it is not feasible to visually represent the learned content in a textual format.
-Thus, future research will need to address the interpretability of these prompts and investigate methodologies that enhance their comprehensibility while maintaining their efficacy.
+In this study, we propose the Semantic Contextual Integration (SCI) network, a novel method that incorporates prompt learning to address cloth-changing person re-identification (CC-ReID).
+The SCI network includes a Semantic Separation Enhancement (SSE) module that generates ID-specific and clothing prompts to effectively isolate clothing semantics (negative factors) from ID information, thereby preserving high-dimensional ID cues (positive factors) in visual features.
+Additionally, the Semantic-Guided Interaction Module (SIM) leverages these textual cues to guide and enhance the discrimination of visual features, facilitating seamless interaction between the visual and textual branches.
+Our work provides valuable insights and suggests promising directions for future research on the role of prompt learning in CC-ReID.
+Although learnable prompts outperform the manually crafted prompts used in CLIP, they often lack semantic clarity and remain uninterpretable to human observers.
+Consequently, the learned prompts cannot be directly visualized in a human-readable format.
+Future research should address the interpretability of these prompts and develop techniques to enhance their comprehensibility without compromising their discriminative power.